OpenStudy (anonymous):

unsimplify 3 radical 2

OpenStudy (bahrom7893):

radical 18

OpenStudy (bahrom7893):

You just asked this.. 3sqrt(2) = sqrt(3^2 * 2) = sqrt(9*2) = sqrt(18)